Show that the space L[a,b] of all square integrable functions on the interval [a,b] is a linear space over a vector field R


1
Expert's answer
2021-11-04T09:16:37-0400

A function y(x) is said to be square integrable if

"\displaystyle{\int^{\infin}_{-\infin}}|f(x)|^2dx<\infin"


for linear space:

"(f+g)(x)=f(x)+g(x)"

"f(ax)=af(x)"


for the space L[a,b] of all square integrable functions:

if

"\displaystyle{\int^{\infin}_{-\infin}}|(f+g)(x)|^2dx<\infin"

then

"\displaystyle{\int^{\infin}_{-\infin}}|f(x)|^2dx+\displaystyle{\int^{\infin}_{-\infin}}|g(x)|^2dx<\infin"


"\int f(ax)dx=a\int f(x)dx"


so, this is a linear space over a vector field R